### v2

Abstract
We present the results of an experimental investigation into buoyant rigid spheres rising through highly concentrated collections of hydrated hydrogel particles. The volume fraction of particles is such that the mechanical properties of the material are intermediate between a very viscous fluid and a soft solid. Despite the established time dependent, non-Newtonian character of hydrogels, we find that when the surface of the material is free, an immersed buoyant sphere rises with a constant speed. The effects of the motion are observed to be highly localized around the sphere. When the stress exerted on the material is changed by varying the mass of the sphere, its terminal velocity is found to depend exponentially on its buoyancy. Qualitatively distinct behavior is found when a solid lid is placed on the surface of the material. In this case, a seemingly thixotropic, sublinear time-dependent motion is found. It is observed that linear motion of the sphere is accompanied by flow at the surface of the material whereas fluid movement is suppressed when a lid is present. We use these observations to provide a hypothesis which links the exponential stress dependence of the rheology of the material to the effects of the boundary conditions on the kinematics of the intruder.
Contents
I Introduction
We report the results of an experimental investigation into the slow motion of rigid spheres through a soft solid. The soft material comprises a highly concentrated collection of hydrated hydrogels. The particle concentration is such that the swollen particles form an interconnected network of highly elastic, porous material with no visible regions of water in the interior or at the surface of the sample. The buoyancy driven motion of the sphere involves complex interactions determined by viscoelastic and capillary forces so that the material response is a mixture of fluid-like deformation and solid-like resistance. Applications of such behaviours are to be found in soft robotics and soil sciences [1].
By way of analogy we make connections with the falling bob viscometer which is a standard laboratory method for measuring the viscosity of very viscous fluids. The method involves applying Stokes law for the motion past a sphere [2] to the terminal speed of rigid particles falling or rising in very viscous fluids or at very small lengthscales. Stokes law applies when where the Reynolds number with , the terminal speed of the sphere of radius and is the kinematic viscosity of the fluid. In practice, the law is valid for many viscous oils and fluids such as glycerine and syrups. The falling bob viscometer technique is nowadays highly refined [3, 4] where effects from the walls of the container [5, 6] or the finite length of the apparatus [7] are taken into account. Moreover, this configuration has been used to investigate the motion of rigid spheres through suspensions of neutrally buoyant spheres over a range of volume fractions [8, 9] and suspensions of fine rods [10]. The method has also been extended to falling cylinders in narrow tubes [11].
Typically, it is assumed that the fluid properties are independent of time and speed for falling-bob viscometers to be used effectively. Moreover, in the case of suspensions, there is a tacit assumption that the kinematics of the particles in the viscous fluid is a direct metric of the local Newtonian rheology of the material. In this case, the local shear stress is proportional to the local shear rate via the viscosity. However, falling bob viscometers can also be used to provide deeper insights into the flow behavior of non-Newtonian fluids, such as viscoelastic materials [12, 13]. For example, in a study of a shear thinning fluid it is found that the distance the sphere travels is [14]. An explanation for this non-linear behaviour is provided by [15] who postulate that a build up of stagnant fluid around the sphere will lead to an increase in its effective size over time and hence Stokes drag increases accordingly. In shear thickening fluids, a similar coupling has been observed between the effective size of the intruder and the viscous response of the material surrounding it for the non-linear slowdown of an intruder [16], which may lead to oscillatory velocity fluctuations [17]. These examples highlight that for non-Newtonian fluids, the observed intruder kinematics can only be reliably interpreted when a detailed knowledge of the constitutive relationship of the fluid is used in combination with an appropriate model of the boundary conditions.
The focus of our current study is on a class of non-Newtonian materials, namely mixtures of hydrated hydrogel particles. Indicative of the ambiguous mechanical nature of these materials, we also refer to them as hydrogel suspensions when we discuss their overall fluid-like properties, or hydrogel packing when we discuss the static properties of the sedimented hydrogel particles. They have many remarkable features, including sensitivity to the frictional properties of the hydrogels, non-Newtonian flow profiles, critical packing fraction behavior in their high flow rate rheology and various forms of creep behavior [18, 19, 20, 21, 22]. Control of the properties of hydrogel suspensions in water can be achieved via the friction of hydrogel particles [18], packing density [23, 24, 20, 21] and even using the valency of added ions [25]. In previous studies, the non-Newtonian fluid properties of dense hydrogel suspensions is known to be paramount. For example, hydrogel suspensions have been shown to behave as Herschel-Bulkley fluids [18] when probed with standard rheological methods under steady state and constant driving rate conditions.
I.1 Constant speed motion
Our primary goal is to use the conceptually simple geometry of a falling-bob viscometer to investigate our counterintuitive observation that under some circumstances a buoyant, rising sphere can travel at a constant speed through samples of hydrated hydrogel suspensions, despite their apparent thixotropic properties. Results are obtained over a wide variety of sample conditions and sphere sizes and are surprising since hydrogel suspensions are materials with well-estabished non-Newtonian properties. Indeed, our previous work revealed a creep flow motion in which the speed typically decreases with the square root of time [21, 22], indicating the presence of significant time dependent, non-Newtonian behavior at these constant stress driving conditions. Hence, observation of constant speed motion that depends exponentially on the driving stress over a wide parameter range is a novel and surprising result which reinforces the need for establishing a connection with the constitutive equations that govern soft sphere suspensions.
In our investigations, we systematically vary the boundary conditions of the hydrogel particle mixture to find that the difference between the constant speed motion observed in the current work and the nonlinear creep motion observed previously originates in the boundary conditions applied to the material. Despite the rheology of hydrogel suspensions being an intrinsic property of the microscopics of the material, we establish that its local kinematics can be affected by the boundary conditions, which is highly unusual. Our work suggests that further theoretical work is required to describe how the constitutive behaviour of hydrogel particle suspensions is affected by the presence of specific boundary conditions. Below, we offer a hypothesis that connects the exponential stress sensitivity in the rheology to the change in thixotropic behavior under different boundary conditions.
I.2 Dependence on material properties and boundary conditions
The motion of rigid intruders is described using concepts based on the observed constant terminal speed of the spheres. We confirm that the material flows in a continuous way and the intruder moves smoothly as resolved using high speed video imaging. We use these observational facts to assign an effective constant viscosity at a prescribed value of the stress, even though it is clear that the material under investigation is non-Newtonian. In fact, the effective viscosity characterization of the material reveals an exponential stress dependence of this property on the buoyancy stress exerted by the intruder. This suggests Arrhenius or Eyring-type [26] mechanisms to be a leading factor in the dissipative mechanisms of the material, albeit that they occur at much lower stress levels than the traditional pressure sensitivity of ordinary viscous fluids [27]. Even so, the exponential stress dependence is also observed in the non-Newtonian, time-dependent creep behavior previously observed in hydrogel suspensions [21, 22].
The dependence on temperature of the observed suspension flow behavior is observed to be significantly greater than the temperature dependence of water alone. This finding indicates that the effective viscosity of the suspension is not dominated by the solvent viscosity alone e.g. via a local mechanism such as lubrication layers between the particles, as these would scale sinking or rising speeds in proportion with the water viscosity. In fact, particle image velocimetry analyses reveal that the deformation field of the hydrogel suspension is highly localised around the intruder.
Despite the localization of the flow, we show that the constant speed motion is consistently changed to a sub-linear time dependence of the displacement when a buoyant sphere rises towards a rigid lid floating on the upper boundary of the hydrogel suspension. We therefore systematically explore different boundary conditions of the hydrogel suspension to find that constant speed motion is found for all intruder motions towards a free boundary, including a boundary formed with a liquid on which the hydrogel suspension is floating.
II Experimental set-up
A front view of the apparatus used in the majority of the experiments is shown in Fig.1. It comprises a mm o.d., mm long plexiglass cylinder which has a mm thick wall. The cylinder is aligned vertically in the laboratory frame of reference and is mounted on a mm thick aluminium base which has a machined mm diameter circular groove containing an ’O’ ring. A silicone sealant is used to attach the cylinder to the base. The cylinder contains two litres of triply boiled Oxford tap water. Boiling is used to reduce the amount of dissolved gases in the water (see for example [28]) and hence help prevent the formation of bubbles over the periods of the experiment, which were typically months. It is observed that the presence of bubbles can significantly affect the reproducibility of the experimental results, probably as a result of the effect of compressibility and surface tension on the packing mechanics [21] and the relatively high Laplace pressure in small bubbles in comparison to the other stress scales that apply in the material. Repeatable results are obtained over the entire period of the experiment using this sample preparation protocol.
II.1 Upper Boundary Conditions
The cylinder is sealed at the top by a tight-fitting lid containing a neoprene ’O’ ring. The sealed container helps to minimise any evaporation of the water. There is a cm gap between the upper surface of the sample and the underside of the external lid. The upper boundary of the sample is a free surface for of the experiments and a printed plastic lid is placed on the surface of the sample in the remainder. The lid is 3D printed in plastic (PLA), hence it is partially aerated and weighs g. Its bottom surface is smooth since it is formed in contact with the heating element of the 3D printer, and is hence comparatively free of defects. It is placed on the surface of the hydrogel packing and a thin layer of water is squeezed on onto the upper surface. The lid is lightweight and the results of an investigation into the effects of confinement [22] indicate its contribution to the bulk stress on the material is minimal.
Finally, a third upper boundary condition is imposed using a steel mesh placed on the surface of the sample. The stainless steel mesh was circular and contained an array of approximately square holes formed by mm wires which cross orthogonally. Meshes of this type are used in devices such as cafetieres. It has a diameter of mm, thickness mm and a pore size of mm and weighs g. Hence water may pass through the mesh easily, while hydrogels are held in position by the mesh. The surface of the mesh is very rough, providing a no-slip boundary condition to the hydrogels.
A time lapse set of images are superposed on the frontview of the apparatus in Fig. 1 to give an impression of the upward motion of a ping-pong ball. To initiate the experiment a buoyant ping pong ball is pushed down to the centre of the base of the cylinder using a mm diameter metal rod which has a hoop at one end. Smaller spheres were manipulated with a similarly shaped mm diameter rod. A settling period of minutes is allowed before the ball is released on the centre line and the hoop is carefully moved to the side of the cylinder. The timescale for the insertion and release of the sphere are typically much shorter than the transit time through the sample.
II.2 Particle tracking
The motion of the sphere is tracked by capturing images using a computer-controlled camera at set intervals ranging from to seconds depending on the buoyancy of the ball. In the case shown in Fig.1 the images are taken at second intervals using a Nikon digital camera. The recorded images are stored on a PC using the software DigiCam and are processed offline using the software ImageJ. The speed is estimated from the resulting distance versus time plots.
II.3 Sample preparation
Sample preparation proceeds as follows. Measured amounts of dry hydrogel powder are slowly mixed with the water which is gently stirred and left to settle for two hours between additions. The particles swell as they absorb water and sink to the bottom of the cylinder as they are slightly more dense than the fluid. The sample is gently stirred between each charge. The hydrogels absorb the water efficiently and form a layer of swollen particles at the bottom of the cylinder with a clear layer of fluid above. Measurements of the depth of the hydrogel layer as a function of the added mass of dry powder is linear as in previous experiments [21]. This finding is consistent with the hydrogel particles being essentially frictionless spheres [29, 18, 30], which settle to a random close packing. The height of hydrogel particle interface reaches the surface of the water at a hydrogel concentration g/L where is the weight of the dry powder divided by the total volume of water in grams per Litre (g/L). Once the layer has reached the surface, the addition of further quantities of dry hydrogels does not give any detectable change in the layer height and presumably there is a small compression of the particles throughout the layer. The average size of the swollen particles is measured to be mm.
The value of used in the majority of the experiments reported here is set to g/L. We choose to use the concentration to set the properties of the sample. As mentioned above, this is a robust control parameter which produces repeatable experimental results. The more common parameter of volume fraction used in studies of suspensions of hard spheres [31] is difficult to estimate with soft materials. Recently, it has been shown that accurate estimates of the volume fraction of hydrogels in suspension [32] can be made only with specialized imaging equipment, such as X-ray facilities.
The majority of the experiments were performed over a period of approximately one year in a laboratory where the measured temperature of the sample is °C. Over some periods the temperature varied outside of this range and the effects of this are reported below. However, the majority of the data given below is collected with the temperature within the given range. We will see below that this temperature window gives at most a variation of measured travelling speeds, which is insignificant for the effects considered, given the large range of travelling speeds explored via the variation of buoyancy stress and intruder size.
Experiments were also performed at the University of Amsterdam, under standard laboratory temperature control. In these cases, a suspension of hydrogel spheres was formed in a square plexiglass box of dimensions cm following our established protocol using dry hydrogel spheres with triply boiled Amsterdam tap water at a concentration of g/L unless otherwise noted. To initiate the experiments, a rod with a hoop was mainly used to place the sphere at the bottom, and in other cases a small magnet was placed inside the ball. An external magnet was used to hold and release the ping pong ball from the bottom of the container.
In the Amsterdam experiments the software TroublePix was used with a Lumenera Lt225C high speed camera to record a video of the motion at a frame rate of frames/second. Still images were captured using a Nikon D5200 camera with ball tracking imaging performed using MaxTraq and Matlab. The flow field was measured using Particle Image Velocimetry (PIV). In this, the plane of motion of the sphere was illuminated by two vertical green laser sheets. Spatial averaging which co-moves with the intruder was used. Matlab integrated software “PIVlab” was employed to analyse the sequences of images.
II.4 Intruder specifications
Ping pong balls are manufactured from Acrylonitrile Butadiene Styrene (ABS) plastic, have a diameter of mm and weigh g. Their size, sphericity and weight are strictly controlled by the International Table Tennis Federation [33] and the dimensions of each ball are checked using a vernier gauge. The diameter ratio of the sphere to container where is the diameter of the sphere and is that of the cylinder. Changing the buoyancy of a ping-pong ball is relatively straightforward as water can be injected into the ball using a hypodermic needle, the injection hole sealed and the weight of the ball measured using an electronic balance. The buoyancy force can then be estimated using the difference in mass between the sphere and the displaced water. We assume that the stress that an intruder exerts on the hydrogel suspension is its effective weight divided by its projected cross sectional area. The main set of experiments are performed using ping-pong balls with a hydrogel sample with g/L. A second set of experiments were carried out using a hollow mm diameter polypropylene ball (Precision plastic ball co.). Injecting water and weighing is again used to change and estimate the buoyancy. The value of is g/L. A third set of experiments are performed with hollow balls made from a range of plastics (polyamide, polypropylene, PTFE, polyethylene) and had diameters mm. These hollow plastic balls were weighed using an electronic balance and the the weight was used to compute the buoyancy stress. The value of used was g/L for these experiments. Finally, in one experiment, a glass sphere with a diameter of mm was used.
III Overall rising intruder dynamics
The results of a typical set of experiments are shown in Fig. 2 for a representative data set. It may be seen that the distance a sphere rises in this hydrogel sample is proportional to when there is a free surface at the top of the sample. The inset Fig. 2 confirms the linearity of the displacement dynamics on a double logarithmic scale. The constant speed motion stands in stark contrast with previous square-root time dependence observations [21, 22]. From the systematic exploration of intruder motion reported below, we deduce that the (end) boundary plays a defining role in the timescale of the motion of the sphere through the sample. We first present a systematic investigation into the constant speed motion of buoyant spheres moving towards a free boundary.
III.1 Time Histories with a Free Surface
As discussed above, the buoyancy of the ping-pong ball can readily be changed by adding small measured quantities of water to its interior. The time histories obtained over a range of buoyancies are shown in Fig. 3 (a). It can be seen that a straight line provides a satisfactory least squares fit to each time history with typically a 3% root mean squared error, and hence there is a linear dependence of distance on time. A wide range of buoyancies was explored using this approach, from a sphere without added water (Pa) and the total rise time is minutes to one with Pa, where the sphere takes hours to rise through the sample. We show all the behaviour for the experiments with ping-pong balls in a sample with g/L and a free upper surface in Fig. 3 (a).
A fixed terminal velocity is found for each value of in the specified sample. This suggests an interpretation of the results using the laws of standard Stokes viscometry might be possible. Of course, the connection cannot be complete since we shall see that the rise velocity depends exponentially on buoyancy; the hydrogel sample has clear non-Newtonian properties. Even so, models of rigid particles in suspensions [12, 34] suggest that an effective viscosity can be ascribed to suspensions of particles in fluids. Following this approach we extract an effective viscosity of the hydrogel medium from the constant intruder velocity data by converting the directly measured terminal velocity of the rising sphere into a value with the units of Pas, using the traditional Stokes variables as the radius of the sphere, is the difference in density between the intruder and hydrogel medium respectively and for gravity:
In this interpretation, is the effective dynamic viscosity of the fluid. Strictly speaking, this formula is only valid for a Stokes flow in an unbounded domain. If the fluid medium used here were a Newtonian viscous fluid, wall and side-wall effects would be significant for the geometry used here. However, as was found in [35] and reported below, motion in the hydrogel suspension is only significant close to the sphere, so we choose to use the simplified analysis for the present study. We will show PIV measurements below, which also indicate that the flow field is localised around the sphere.
The effective viscosity approach enables us to rescale all the displacement data acquired for one sample onto a master curve that follows . We show the result of this analysis in Fig. 3 (b). An excellent collapse over three orders of magnitude confirms the constant speed dynamics of an intruder rising towards a free surface.
The collapse of the data provides an important insight. It is possible that the time dependence observed previously [21, 22] could arise from a gradient in the properties of the sample resulting from concentration variations or the effects of hydrostatic pressure, for example, or other gradients because of the sensitivity of the viscosity to packing density around the random close packing limit [23]. The constant speed motion, instead, highlights that concentrated samples of hydrated hydrogels which do not display spatiotemporal density variation effects can be consistently produced.
A graph of the effective viscosities of the time histories presented in Fig. 3 (a), (b) versus buoyancy stress is given in Fig. 3 (c). The dependence is exponential as shown by the linear fit to the data since the ordinate has a logarithmic scale. Hence, with a constant that depends on the hydrogel density, among other factors. This exponential dependence of propagation speeds also depends on packing fraction [21] and global confinement stress [22], which is consistent with material response previously reported for these materials.
III.2 Sample robustness
We show the constant speed motion in the time histories is also consistently observed for three different values of and g/L in Fig. 4 (a). For these samples, there are good least squares fits of a straight lines with the same accuracy as reported above. This indicates a constant velocity approximation is an appropriate descriptor for the dynamics of the sphere rising towards a free surface. The terminal speed of the sphere can be estimated from the slope of the least squares fitted line through and an estimate of Fig. 4 (a) is made in each case. The estimated value of can be used to rescale the time history plots and a convincing collapse of the data is found as shown over four orders of magnitude. Each measurement indicates an exponential dependence of on . This trend is shown in Fig. 4(b) plotted as a function of buoyancy stress . As expected, based on previous work [22], for each sample, the characteristic stress of the is slightly different.
We conclude that despite the constant speed motion of the sphere, the fluid medium is not Newtonian: there is an exponential drop in effective viscosity as the buoyancy is increased. Specifically, the log-linear plot of the slopes of the versus plots shown in Fig 4 indicates that the propagation speed of the sphere depends exponentially on the buoyancy stress with stress factor , which varies between to Pa. This finding is consistent with previous results for stress dependence of creep at both local [21] and global [22] levels for both spheres and cylinders.
III.3 Results with smaller diameter spheres.
Smaller diameter buoyant spheres are used to test the robustness of the results. These include a hollow plastic sphere of mm in diameter, which is filled with different amounts of water to reach the desired weight. Hollow PTFE, delrin and polypropylene spheres of mm, mm and mm are also employed. These intruders are used with hydrogel samples at and g/L. We find that the rising dynamics for the smaller diameters are in agreement with the data for the ping-pong balls over the entire range of buoyant stresses investigated as shown in Fig. 5. We conclude from these results that finite size effects of the spheres are not significant.
IV Flow structure within the hydrogel suspension
The constant speed motion of the rising intruder begs the question if the flow field around the slowly moving sphere bears similarity to a Stokes flow. To investigate this, we perform experiments with ping pong balls rising towards a free surface in a hydrogel suspension in a square container illuminated by a pair of green laser sheets as shown schematically in Fig. 6. We analyze the measured flow fields around the sphere using the software PIVlab [36] applied to the videos recorded with the high-speed camera. The videos were pre-processed using intensity capping to reduce the noise from the brightest spots which arise from scattering of the laser illumination by impurities in the hydrogel suspension. We also remove the image of the ping pong ball from the region of interest for the PIV analysis using a mask.
The flow field analysis is performed using a Multipass FFT window deformation algorithm. In particular we used three passes with an interrogation area of 80x80, 40x40 and 20x20 pixels respectively (corresponding to approximately 18x18, 9x9 and 5x5 mm) with 50% overlap and with the “Extreme” correlation robustness option. This option has been shown to yield results with high signal-to-noise ratio in the presence of noise [37]. We also employ a standard approach of using different frame jumps to measure the slow velocities far away from the moving intruder [38].
Finally, we remove the few outliers (which is a result of out-of-plane motion or poorly lit areas) through two velocity validation algorithms, automatically implemented in PIVlab. We first filter the velocity vectors based on the magnitude of each component, requiring that it does not deviate by more than 8 standard deviations from the mean, and then we apply the local median filter [39]. This procedure yields a Valid Detection Probability (VDP) of higher than 98%. A typical PIV result from an experimental video is available as Supplementary Information.
The resulting PIV fields are averaged within a domain that co-moves with the steadily rising sphere obtained over 700 frames. The velocity field representing left-right motion is shown in Fig. 6 (b). The quadrupolar flow structure as required by mass conservation can be clearly distinguished, and there is a weak vertical asymmetry in the profile. The velocity component parallel to gravity is shown in Fig. 6 (c). The colour scheme highlights that the upward speed of the sphere is significantly higher than the return downward speed of the water in the hydrogels. This is reasonable since the downward counterflow required by conservation of mass is spread over a wide area surrounding the sphere. To highlight the structure of the additional flow field induced by the shear between the sphere and the hydrogel sample, we constrain the colour scale in Fig. 6 (d). There is a significant wake flow dragged upwards with the sphere, and there is also a large volume of hydrogel particles pushed ahead of the sphere. However, adjacent to the sphere, the hydrogels are mainly stationary, even though a no-slip boundary condition exists between the water and the sphere. Our measurements are accurate up to a distance of mm (approximately the size of a particle) from the spherical intruder. Combining the velocity components into permits a comparison of PIV results with tracking measurements and the equatorial velocity component with the Stokes solution, Fig. 6 (e). We observe that the hydrogel fluid velocity profile has a very short decay length when compared with the long range Stokes solution.
Implicit in our arguments thus far is that the medium flows in a continuous way. In practice our sample comprises a mixture of polydisperse, deformable, porous, hydrogels and water which could give rise to discrete displacement steps from e.g. stick-slip or other intermittent dynamics. We are unable to detect intermittency when we use higher spatiotemporal resolution. In Fig. 7 the results of an experiment on the motion of a ping-pong ball recorded using the high speed imaging system described above show there is no evidence for discrete steps in the motion. Note that these results were obtained using the square container setup. The same qualitative behaviour was observed in separate experiments with a circular container, indicating that side-wall effects on the motion are also not significant.
V Temperature and Diffusion effects
The constant speed behaviour suggests that a water-based lubrication mechanism may play a role in the dynamics of the rising sphere. We can test this hypothesis by measuring the effective viscosity of a sample under varying environmental temperature. The sample temperature is recorded in all experiments and the buoyancy stress is kept fixed for this particular investigation. We plot in Fig. 8 the effective viscosity of a hydrogel sample versus temperature with g/L with a buoyancy stress of Pa. The temperature range covered is to °C. Also included in Fig.8 are data points for water taken from tabulated values [40]. The data sets are normalised to be unity at °C. The fitted lines have an Arrhenius form; it can be seen that the hydrogel sample has a significantly stronger dependence on temperature than water. We conclude that a simple water-based lubrication mechanism, for example in the observed slip between the sphere and the hydrogel sample found here and reported elsewhere [35], cannot be the only factor that plays a role in the effective viscosity of the two-phase medium. We speculate that other temperature dependent mechanisms contribute to the effective viscosity of the hydrogel sample. One of these could be the temperature dependence of the stiffness of the hydrogels [41], which controls their poroelastic behavior [42]. However, poroelastic effects typically have a timescale that is the ratio of the stiffness and the viscosity, which for the system observed here both decrease with temperature, partially compensating this effect.
Another local mechanism that can drive motion of an intruder is local particle diffusion in the hydrogel suspension. It is thus interesting to understand how fast the intruder moves with respect to diffusive processes in the hydrogel suspension. The intruder motion is of the order of to m/s. There are two types of diffusive processes involved. At the molecular level motion within the water is one, but also the hydrogel particles close to the sphere are affected by passage of the intruder [43]. More intensive Particle Tracking Velocimetry could aid in extracting the mean squared displacment of the hydrogel particles. With the intruder speed and a characteristic lengthscale , one can estimate a Peclet number . From the available dyed tracer particle measurement we estimate that hydrogel particles do not diffuse more than a millimeter during the course of an experiment, putting an upper bound on their mean squared displacement m2/s. The diffusion constant of water is around m2/s, suggesting that the Peclet number should be larger than one for both the linear and sublinear scenarios. This simplified interpretation may be more subtle in practice: for example we expect to be dependent, as the diffusive motion of the hydrogels is driven not by thermal motion of the solvent molecules, but by the movement of the intruder [43], involving a self-fluidization process that is commonly observed to be present for dry granular materials [44, 45, 46]. The typical lengthscale involved is also not easily identified. We thus see several avenues for future research to explore the role of microscopic mechanisms that may contribute to driving the slow flow in hydrogel suspensions.
VI End boundary conditions
All the data we have discussed thus far displays a linear dependence of distance on time for spheres moving towards a free boundary. As discussed in the introduction, previous results for the motion of heavy spheres through hydrated hydrogels [21, 22] which correspond to cumulative 200+ measurements show . The main difference between the two configurations is that in the present study, the ping-pong balls rise towards a free surface whereas the heavy spheres and cylinders used in earlier work were descending towards a rigid boundary. A question which may be asked is whether the approached end boundary condition plays a role in the motion of the intruder.
We therefore compare the example intruder motion data from Fig. 1 obtained from an intruder moving towards a free boundary, with observations of the same intruder motion moving in the same hydrogel sample towards a light rigid lid floating on the hydrogel sample. We observe that the motion of the buoyant sphere moving towards the floating lid is clearly sub-linear. The example data for sub-linear motion shown in Fig. 9 is best fitted with a power law of with , as can be seen in the double logarithmic inset in Fig. 9 .
After systematically exploring this positive buoyancy case for a hydrogel suspension confined between two solid boundaries (a lid at the hydrogel surface and the container wall at the bottom), we also investigate two other boundary conditions. Firstly, we measure the motion of an intruder with a positive buoyancy and a packing confined by a stiff, porous metal mesh. Secondly, we investigate the motion of intruders with both positive and negative buoyancies with a hydrogel sample confined between two fluid boundaries (air at the top, heavy liquid at the bottom). Finally, we consider flow field aspects that pertain to the boundary conditions, visualizing the strength of the displacement field close to the boundary at the top of an unconstrained hydrogel packing. These additional observations force us to conclude that the end boundary conditions play a crucial role in setting the overall flow behaviour of intruders in the hydrogel packing.
VI.1 Sphere moving towards a rigid floating lid.
To systematically address the issue of the effects of end boundary conditions, a sequence of experiments are performed where a printed plastic lid sits on top of the swollen hydrogels. The sample used as before has g/L and the buoyancy of the sphere is adjusted by adding or subtracting water to and from the interior of the ping-pong ball as described previously.
As outlined above in the discussion of the results in Fig. 1, the qualitative difference in the time series plots for cases with and without a lid is clear. At the same value of and buoyancy, the sphere rises with a linear distance-time relationship with a free surface . However, it is sub-linear, approximating previously observed scaling, when the lid is present. Note that a dependence of distance travelled on is qualitatively different to the end effects in falling-bob experiments on viscous Newtonian flows [7].
The sublinear dynamics observed suggests a systematic exploratiom of the effects of a lid on behavior is required. The results of tests with ping pong balls of various buoyancies are shown in Fig. 10 (a). As shown in the example data from Fig. 9, we observe a clear sublinear time-dependent behavior of the displacement of the sphere. In order to extract the time dependence of the displacement of the rising sphere, we again fit to each data set
Here, is a prefactor set by the speed of motion. Despite the non-Newtonian nature of the flow behavior and the unusual dimensions of , we refer to it as the effective viscosity. We verify which rescaling coefficient provides the best collapse of the data; plotting results in a linear time dependence. We find that provides the best collapse of the data and show this result in Fig. 10b. We verify the exponential dependence of prefactor that we use to characterise the strength of the sublinear behavior in Fig. 10c. Since the same hydrogel sample was used here, we can compare the exponential dependence on the buoyancy stress for the sublinear case with the case of the sphere rising towards a free surface as discussed in Sec. III.1 (dash-dotted line). The data for the sublinear case clearly follows the same scaling of , with the same Pa as used for the free surface case. We cannot expect a perfect match in the functional form, as absorbs some of the physical prefactors that we could separate out in the constant velocity case. We conclude that motion of a sphere towards a rigid boundary is sublinear by way of contrast with the motion towards a free surface, which is linear.
VI.2 Sphere moving towards a rigid base — comparison with previous work
In previous work [21, 22] we performed experiments on sinking intruders that moved towards the rigid base of the container. These additional data enable us to compare the observed time dependence with the cases discussed in this manuscript. We extract the data from the open repositories, e.g. [47] and perform the same fitting procedure to estimate the behavior for these data sets. To allow for a systematic comparison of our results, we publish our kinetic data and analysis scripts in an open repository [48].
We can extract the exponent for which a power law model provides the best fit. We note that there is no canonical definition of a “best” fit, as fit quality as measured by a reduced depends also on weights associated with data, and even with an optimal , systematic deviations in the residual of the fit are neglected. Our choice for “best” is hence simply the best goodness of fit estimate as returned by the MATLAB-based fitting routine.
Applying this same standard across all data sets from two previous publications and the current one, yields the results as shown in Fig. 11, with over 300 sets fitted. The statistics of the distribtion of for each case of “rigid” boundary, “floating lid” and “free” surface is indicated with a violin plot, including a separate scatter plot for the mean and standard deviation of the distribution (error bar). We observe a clear clustering of exponents around 0.5 for all the previously published data, in which an intruder was sinking towards a rigid base. In the current study, some data was obtained with an intruder rising towards a rigid lid floating on the hydrogel packing; the exponents for these experiments center around 0.7. The exponents for the creeping flow measurements for intruders that move towards a free boundary clearly center around , confirming again the strong influence of the boundary on the flow behavior of the hydrogel suspension. Lastly, for the data sets of the previous investigations we can obtain a satisfactory collapse as obtained for the current data as shown in Fig. 10 (b). The temporal resolution in our previous work was lower than used here, but the exponent which provides a satisfactory collapse for these datasets is for data from [21, 22] albeit with increased scatter. All data and analysis scripts for all our work can be found in open data repositories [47].
VI.3 Rising towards a surface with a mesh
We show in Fig. 12 four time histories with hydrogel density g/L. Two of the time histories were measured with a free surface and two with a metal mesh which sits mm below the surface of the packing. The mesh prevents the hydrogel particles at the surface from moving during the transit of the ping-pong ball, yet it is sufficiently porous that water will flow relatively unimpeded. This boundary condition is rough and hence suppresses vertical and radial motion of the hydrogels at the top of the packing. However, the mesh is also very light. The results indicate that the mesh has very little effect on the constant speed motion of the ping-pong ball. Moreover, the surface of the sample remains flat during the motion of the ball, indicating that the constant speed motion is not an end effect imposed by having a surface structure imposed, suggesting that flexible rough lids would also not create sublinear motion. We cannot exclude the possibility that a smooth yet flexible boundary such as the surface tension boundary conditions from [21] can nevertheless be essential to create sublinear motion.
VI.4 Sinking towards a free, fluid boundary.
Experiments were performed to investigate whether the constant speed behavior exists for heavy spheres that sink towards a “free” boundary formed under the hydrogel suspension by placing a heavy immiscible fluid at the bottom of the container. In these experiments, the hydrogel suspension or was created on top of a cm liquid layer of Galden 135 fluid. The hydrogel suspension was made with deionised water, in which the hydrogel particles swell more easily. This accounts for the lower value of used here in comparison to the typical values used in the ping-pong ball experiments. We confirmed that this lower value of produced a sample in which the entire water volume is densely packed with hydrogel spheres.
A water-filled ping pong ball is used in the initial experiments. Hence, it is negatively buoyant in the hydrogel suspension, but positively buoyant in the heavy Galden liquid. The motion of the filled ping-pong ball achieves a constant velocity while fully immersed in the hydrogel suspension but slows to a standstill at the interface of the hydrogel suspension and the Galden. The observations are found to be repeatable. The second set of experiments involves a glass sphere (kgm-3), which is negatively buoyant in both the hydrogel suspension and Galden. The hydrogel suspension used in the second experiment is slightly denser to set the sinking time to be consistent with previous observations. Despite the differences between the two sinking experiments, the sphere reaches constant velocity as indicated in Fig. 13 (a). Rescaling the displacement data with the effective viscosity of the respective suspension-intruder pairs collapses the data in a straight line with slope , confirming once more that even if the motion is in the direction of gravity, an intruder driven by a constant stress moving towards a free boundary of a hydrogel suspension maintains a constant speed.
VI.5 Rising towards a free surface: radial displacement field
To investigate the radial velocity profile of the hydrogel suspension during a rising sphere test, the hydrogel suspension is marked with some coloured swollen hydrogel beads, and their motion is imaged from the top of the container. The image shown in the top right panel of Fig. 14 is a composite constructed from still images taken at second intervals during the period in which the sphere is moving towards the surface. Evidence for outward motion from the centre of the ball is clear. Upward movement of the hydrogel surface was also observed; see Fig. 14, bottom right. The flow visualisation observations in Fig. 14 illustrate that the radial flow of the hydrogel suspension above the sphere is substantial, and that boundary of the hydrogel fluid can deform significantly. However, as indicated with the mesh experiments, these deformation fields alone do not seem to be responsible for creating (sub)linear time dependent motion of the intruder.
VII Discussion
Some intuition on the origin of the linear-to-sublinear transition of the intruder dynamics under the change of boundary conditions can be gained by considering a simple viscoelastic model to describe the hydrogel packing as a continuous fluid with an effective viscosity that sets the drag force on an intruder. Let be the effective buoyancy force experienced by the ping pong ball. In the overdamped limit in which we work, a simple force balance leads to the following differential equation
Even for a typical non-Newtonian rheology such as a Herschel-Bulkley fluid, which includes a yield stress, a constant motion perspective is to be expected, if the applied intruder stress is above the yield stress [13]. In this case, would be corrected for the existence of the yield stress and would be affected by the flow index.
In the free boundary case, thus, it seems reasonable to assume a constant viscosity , which leads to a constant velocity. We can then identify the exponential stress dependence of the hydrogel packing [22] as the mechanism to explain the thixotropic fluid response in the presence of a lid. Adding a lid at the top of the packing constrains or confines (at least partially) the hydrogel particles in the volume between the intruder and the lid itself, suppressing their motion and limiting their ability to escape from the aforementioned volume. Our hypothesis is that a modulation of the density in the packing, due to the combined effect of the lid confinement and of the intruder motion, is responsible for an increase in the total dissipation of intruder-induced hydrogel motion close to the rigid boundary. This, consequently, corresponds to a viscosity that grows with time, or, more importantly, that grows as the intruder approaches the rigid boundary, , allowing us to suggest a simple explanation of how the already known (but as yet unexplained) intrinsic rheological properties of the material could produce a time-dependent viscosity, and give rise to a sublinear motion of the intruder. We now develop this explanation step by step.
It was shown in previous studies [22] that the creep rate in hydrogel packings depends exponentially on the intruder stress, , and on the confinement stress acting on the hydrogel particles, . In the current study we confirm the exponential stress dependence in the linear case (cfr. section III) observing . Thus, we expect the dependence on the confinement stress to also hold: .
Our proposed mechanism is that this confinement stress, stored in the contact network of the hydrogel packing, increases as a result of the motion of the intruder towards a rigid boundary, . The intruder, effectively acts as a moving lid, pushing the hydrogel particles along on its trajectory towards the end boundary. The rigidity of the end boundary does not allow the particles to move radially as instead observed in the case of a free boundary (cfr section VI.5), so the internal stresses cannot relax. The hydrogels are, thus, constrained in a volume which cannot deform, the particles are compressed and the local density of the hydrogel packing increases (see Fig. 15 for a depiction of the mechanism). A perfect knowledge of the dependence of the excess confinement stress on the displacement of the ball, , would allow us to solve equation (3), where for small deviation of the confinement stress we could approximate the exponential stress dependence of the effective viscosity to the following
Assuming that , where is an exponent that depends on the end boundary condition and is a proportionality constant, we obtain which describes linear motion for and sublinear motion for . In particular, in the case of a linear dependence of the excess stress on the displacement, , the solution reads
which agrees with observations from previous studies of sinking intruders towards fully rigid boundaries. In the presence of a less constraining boundary condition, such as the floating lid considered in this study, there would be a sublinear dependence of the stress, and consequently of the viscosity on the travelled distance , and therefore the time history of the intruder will be intermediate between the linear free boundary case and the square-root rigid boundary case. Of course, our qualitative model does not substantiate the underlying mechanisms for the exponential stress dependence. We can also not experimentally access the packing fraction of the hydrogel packing, let alone gain clarity on the local variation of such density and their potential role in the viscosity enhancement. However, the quantitative potential of the model is clear.
VIII Conclusions
We performed falling-ball viscometry tests of hydrogel suspensions. We used hollow plastic spheres to observe the dynamics of a rising sphere as a function of time. The spheres rise with in cases with a free surface on the hydrogel suspension. Each run of the experiment has this feature. The ascent rate is found to depend exponentially on the buoyancy stress i.e. and therefore there is clear evidence for non-Newtonian behavior in the hydrogel suspension. We confirm the non-Newtonian behavior with temperature tests and flow field visualization, both of which reveal non-Stokesian dynamics. We contrast the linear rise dynamics with creep flow found earlier to confirm a dependence of the flow behavior on the boundary conditions. An intruder sinking towards a free boundary also shows linear displacement dynamics, and we find that when a freely floating lid is placed on the surface of the material, , with . The qualitative difference between the two types of motion is accompanied by clear evidence of flow at the free surface. We hypothesize a quantitative connection between the change in boundary conditions, the exponential packing stress dependence and the time dependent flow behavior, stimulating theoretical progress while focusing future work on the yet unexplained but consistently observed exponential stress dependence of the flow behavior of hydrogel packings.
Acknowledgments
The authors are grateful to Keith Long for constructing the apparatus used in the majority of the experiments. We also thank several undergraduate students at the University of Amsterdam for their preparation work on some of the rising sphere experiments. Finally, we are grateful to Dominic Vella for his suggestion of using a mesh as the upper boundary and to the Leverhulme Trust who supports TM’s research through an Emeritus Fellowship EM-2024-014/4.